# Unifying Structure and Activation: A Comprehensive Approach of Parameter and Memory Efficient Transfer Learning

### v1

@@ Abstract
Parameter-efficient transfer learning (PETL) aims to reduce the scales of pre-trained models for multiple downstream tasks. However, as the models keep scaling up, the memory footprint of existing PETL methods is not significantly reduced compared to the reduction of learnable parameters. This limitation hinders the practical deployment of PETL methods on memory-constrained devices. To this end, we proposed a new PETL framework, called Structure to Activation (S2A), to reduce the memory footprint of activation during fine-tuning. Specifically, our framework consists of: 1)Activation modules design(i.e. bias, prompt and side modules) in the parametric model structure, which results in a significant reduction of adjustable parameters and activation memory 2) 4-bit quantisation of activations based on their derivatives for non-parametric structures (e.g., nonlinear functions), which maintains accuracy while significantly reducing memory usage. Our S2A method consequently offers a lightweight solution in terms of both parameter and memory footprint. We evaluate S2A with different backbones and conduct extensive experiments on various datasets to evaluate the effectiveness. The results show that our method not only outperforms existing PETL techniques, achieving a fourfold reduction in GPU memory footprint on average, but also shows competitive performance in accuracy with lower tunable parameters. These also demonstrate that our method is highly suitable for practical transfer learning on hardware-constrained devices.
@@ 1 Introduction
Transfer learning has become a prevalent paradigm in computer vision research, where pre-trained models are fine-tuned on target datasets to adapt to specific tasks. [1, 2, 3] With the increasing scale of pre-trained models, the number of parameters has grown significantly, making the full fine-tuning less affordable, especially for recent very large models, e.g., ViT-G [4] and Swin-G [5]. Moreover, a single pre-trained model may be transferred to multiple tasks, and each fine-tuned model requires considerable space to store. To reduce the offline storage cost, Parameter Efficient Transfer Learning (PETL) is proposed to perform partial fine-tuning on either the original or some specialized additional parametric modules. The additional modules are designed to be lightweight and inserted into the original model during fine-tuning, so that we can update the parameters of such small modules while fixing the original parameters. The representative PETL methods include bias tuning [6] (fine-tuning the bias vectors only), LORA [7] (decomposing weight into two low-rank matrices) and Adapter [8] (inserting parameter-efficient modules into the original model) and VPT [9] (appending learnable tokens into ViTs).In practice, the PETL methods are particularly suitable for cloud-based service where we may require a single model to adapt to multiple streaming tasks without changing pre-trained parameters.
@@ 2.2 Neural Network Quantization
Neural Network Quantization(NNQ)[27, 28, 29, 30, 31, 32, 33, 34, 35] has been widely explored as a method of compression and acceleration. In the training-aware quantization, NNQ quantizes the full-precision (32-bit) weight and activation to lower bit-width before forward calculating. In the backward, the gradient of full-precision weight and activation is substituted by the gradient of the quantized counterpart[36, 37]. To discretize the floating-point data, the most direct method is asymmetrical quantization[27]. In this method, the 32-bit data is projected to the discrete set $\{0,1,2,...,2^{N}-1\}$ by the scale and shift factor where $N$ is the bit-width. In addition to quantizing the floating data with uniform intervals, there are many works toward non-uniform quantization. Han et al. [28] quantizes the weights for less storage by clustering the floating data into different clusters. APoT[38] uses additive powers-of-two to represent the floating data. Zhao et al.  [39] quantizes the weights to non-uniform forms by considering prior distribution. Unlike the previous works that utilize quantization to reduce the storage of weights or speed up the computation in inference stage, our primary goal is to reduce the memory footprint of activation in fine-tuning stage. In our work, we quantize the activation of non-parametric layers based on the properties of their derivative, which has not been explored in NNQ.

### v3

# S2A: A Unified Framework for Parameter and Memory Efficient Transfer Learning

@@ Abstract
Parameter-efficient transfer learning (PETL) aims to reduce the scales of pretrained models for multiple downstream tasks. However, as the models keep scaling up, the memory footprint of existing PETL methods is not significantly reduced compared to the reduction of learnable parameters. This limitation hinders the practical deployment of PETL methods on memory-constrained devices. To this end, we proposed a new PETL framework, called Structure to Activation (S2A), to reduce the memory footprint of activation during fine-tuning. Specifically, our framework consists of: 1) Activation modules design(i.e., bias, prompt and side modules) in the parametric model structure, which results in a significant reduction of adjustable parameters and activation memory; 2) 4-bit quantization of activations based on their derivatives for non-parametric structures (e.g., nonlinear functions), which maintains accuracy while significantly reducing memory usage. Our S2A method consequently offers a lightweight solution in terms of both parameters and memory footprint. We evaluated S2A with different backbones and performed extensive experiments on various datasets to evaluate the effectiveness. The results show that our methods not only outperform existing PETL techniques, achieving a fourfold reduction in GPU memory footprint on average, but also shows competitive performance in accuracy with fewer tunable parameters. These demonstrate that our method is highly suitable for practical transfer learning on hardware-constrained devices.
@@ 1 Introduction
Transfer learning has become a prevalent paradigm in computer vision research, where pre-trained models are fine-tuned on target datasets to adapt to specific tasks. [1, 2, 3, 4, 5]With the increasing scale of pre-trained models, the number of parameters has grown significantly, making the full fine-tuning less affordable, especially for recent very large models, e.g., ViT-G [6] and Swin-G [7]. Moreover, a single pre-trained model may be transferred to multiple tasks, and each fine-tuned model requires considerable space to store. To reduce the offline storage cost, Parameter Efficient Transfer Learning (PETL) is proposed to perform partial fine-tuning on either the original or some specialized additional parametric modules. The additional modules are designed to be lightweight and inserted into the original model during fine-tuning, so that we can update the parameters of such small modules while fixing the original parameters. The representative PETL methods include bias tuning [8] (fine-tuning the bias vectors only), LORA [9] (decomposing weight into two low-rank matrices) and Adapter [10] (inserting parameter-efficient modules into the original model) and VPT [11] (appending learnable tokens into ViTs).In practice, the PETL methods are particularly suitable for cloud-based service where we may require a single model to adapt to multiple streaming tasks without changing pre-trained parameters.
@@ 2.2 Neural Network Quantization
Neural Network Quantization(NNQ)[29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] has been widely explored as a method of compression and acceleration. In the training-aware quantization, NNQ quantizes the full-precision (32-bit) weight and activation to lower bit-width before forward calculating. In the backward, the gradient of full-precision weight and activation is substituted by the gradient of the quantized counterpart[41, 42]. To discretize the floating-point data, the most direct method is asymmetrical quantization[29]. In this method, the 32-bit data is projected to the discrete set $\{0,1,2,...,2^{N}-1\}$ by the scale and shift factor where $N$ is the bit-width. In addition to quantizing the floating data with uniform intervals, there are many works toward non-uniform quantization. Han et al. [30] quantizes the weights for less storage by clustering the floating data into different clusters. APoT[43] uses additive powers-of-two to represent the floating data. Zhao et al.  [44] quantizes the weights to non-uniform forms by considering prior distribution. Unlike the previous works that utilize quantization to reduce the storage of weights or speed up the computation in inference stage, our primary goal is to reduce the memory footprint of activation in fine-tuning stage. In our work, we quantize the activation of non-parametric layers based on the properties of their derivative, which has not been explored in NNQ.